# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ 3.2 Main Results
where $\mathbb{P}_{\theta^{*}}$ is the data generating measure for the pair of neighbouring datasets of size $n$, $\bm{X}$ and $\bm{Z}$, with entries generated from the true likelihood $k_{p}(\cdot;\theta^{*})$. In addition, $t_{n}=c_{1}\phi_{n}^{2}/8$, $\psi_{n}$ is defined in Proposition 4, and $c_{5}=\min\{c_{1}\alpha/8,c_{2}\}$. In other words,
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp(\epsilon_{n})\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}$
holds for all large $n$ $\mathbb{P}_{\theta^{*}}$-almost surely.
The takeaway of Theorem 5 is that as $n\rightarrow\infty$, we may choose a sequence of $p_{n}\rightarrow 0$ such that $\epsilon_{n},\delta_{n}\rightarrow 0$ and for $n$ large enough the failure probability falls into the first category of (16), that is $<C_{1}\exp(-nc_{5}\phi_{n}^{2})\rightarrow 0$.
@@
Proposition 2 gives us some clue for how to upperbound the $\epsilon_{n}$ and $\delta_{n}$. Recall that
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\eta_{n}\left[\eta_{n}+\frac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$
@@
Under Assumption 4 and 5, we can prove that $\eta_{n}$ is always finite with $\eta_{n}\rightarrow 1$. We can also characterise its convergence rate (see Lemma 15).
@@
  • In Lemma 13, we prove, under Assumption 2 that
@@
    $\mathbb{P}^{*}_{\theta^{*}}\left(m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})\geq e^{-c_{1}n\phi_{n}^{2}}\right)\leq C^{**}\exp(-c_{2}n\phi_{n}^{2})$
@@
  • The relevant proofs are contained in the Appendix A, apart from the proof for Lemma 9 in the supplement.
Since we assume that $\bm{Z}$ also follows the same data generating measure, $m_{n}(A_{n}^{c},\bm{Z})$ shares the same upper bound as $m_{n}(A_{n}^{c},\bm{X})$. Now, every term in the inequality has a probabilistic upper bound, combining these bounds and the theorem is proved. ∎
Lemmas 10, 12, Proposition 4 are intermediate results such that whenever the Assumptions 1 and 2 hold with respect to uncontaminated likelihood $f(\cdot;\theta)$, the same also hold for the actual contaminated likelihood $k_{p}(\cdot;\theta)$, and in practice, one only need to check the assumptions for the case $p=0$.
@@
Let $p_{n}\rightarrow 0$ be a sequence of positive real numbers. Let $I_{p}(\theta)$ denote the Fisher information matrix for a single data point with respect to the density $k_{p}(\cdot;\theta)$ and $I(\theta)$ denote the Fisher information matrix with respect to $f(\cdot;\theta)$. Then
@@
Thus by picking $p_{n}\rightarrow 0$ such that $\psi_{n}/p_{n}\rightarrow 0$ as $n\rightarrow\infty$, both $\delta_{n}$ and $\epsilon_{n}$ converges to zero. Thus the algorithm is also asymptotically differentially private for any pair of $(\epsilon,\delta)$. Since $p_{n}\rightarrow 0$, the convergence in the Fisher information follows from Theorem 6. ∎
Our results address the case where samples are drawn exactly from the posterior distribution. When the algorithm of interest is a Markov Chain Monte Carlo algorithm (Robert, 1999), i.e., does not output i.i.d. samples, our result indicates the level of differential privacy of such algorithm when it reaches the stationary state. It is important to note that when the Markov chain has converged, the algorithm is irrelevant, and the level of differential privacy only depends on the target distribution.
@@ 3.3 Remarks on Assumptions
The Assumptions 1-3 resembles the set of assumptions in Theorem 8.9 of Ghosal and Van der Vaart (2017). In particular, our Assumption 3 is equivalent to their Assumption (iii). our Assumption 1 is similar to their Assumption (ii) with the semi-metric $d$ replaced by Hellinger distance $h$ and conditioning on a bracketing number instead of covering number. Finally, our Assumption 2 is similar to their Assumption (i) in the sense that both conditions on the prior mass around a shrinking neighbourhood around the true parameter $\theta^{*}$ are sufficiently large.
We adopt the current set of assumptions in order to use the proof techniques of Shen and Wong (1994); Wong and Shen (1995); Shen and Wasserman (2001), which characterizes the rate of convergence the quantity $\mathbb{P}_{n}(A_{n}^{c})$ with a known failing rate, i.e., (16) type bound, which is not available by the Theorems of Ghosal and Van der Vaart (2017).
@@
By Lemma 3, we know the above will hold for most likelihood functions $f(\cdot;\theta)$ when $n$ is large enough. However, due to the constant $c$ being quite small in magnitude, see Lemma B.2 in the supplement, one usually need $n$ to be unrealistically large for Theorem 5 to hold. Thus despite having derived its convergence rate in a non-asymptotic fashion, the result is better deemed as an asymptotic one. We will show in the next section with empirical simulation results that Theorem 5 does give a good intuition as to how $\epsilon_{n}$ and $\delta_{n}$ converge.
Both Assumptions 2 and 3 are conditions on the prior distribution. Let $\bar{S}(r):=\{\theta\in\Omega:\|\theta^{*}-\theta\|_{2}\leq r\}$ denote the $L_{2}$-ball of radius $r$ around $\theta^{*}$. It is possible to show that under mild conditions, fixing a constant $c_{r}$, the set $S_{0,\alpha}(t_{n})\cap\bar{S}(c_{r})$ contains a $L_{2}$-ball with radius proportional to $t_{n}$.
@@
Let $S_{0,\alpha}(t_{n}):=\{\theta:\rho_{0,\alpha}(\theta^{*},\theta)\leq t_{n}\}$ defined as in Definition 2. Suppose that there exists $\alpha\in(0,1]$, and $c_{r}<\infty$ such that
$\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\int\left[\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)}\right]^{\alpha+1}\|\tilde{\bm{\nabla}}_{\theta}f(\bm{x};\tilde{\theta})\|_{2}\mathtt{d}x<M_{c_{r}}$
where $M_{c_{r}}$ is a constant that depends on $c_{r}$, and $\tilde{\bm{\nabla}}$ is supremum directional derivative defined in Definition 7. Then $\bar{S}(t_{n}/M_{c_{r}})\subseteq S_{0,\alpha}(t_{n})$.
@@
The condition holds for most likelihood functions, since ~$\|\tilde{\bm{\nabla}}_{\theta}f(x;\theta)\|$ is usually integrable already. If $\theta$ does not control the dispersion, then $f(x;\theta^{*})$ and $f(\bm{x};\theta)$ have matching rates of decay and hence the expression is integrable. Otherwise, one usually can choose $c_{r}$ small enough such that the expression remains integrable.
@@
One may use a similar argument to show that the Hellinger distance and the $L_{2}$ distance are locally related (See Appendix C) so that $\phi_{n}$ and $\psi_{n}$ can be chosen to decay at the same rate. Then, the rate of $p_{n}$ can be chosen freely as long as $\psi_{n}/p_{n}\rightarrow 0$.
Note that we may choose $\Omega_{n}$ in Assumption 3 as an expanding $L_{2}$-ball. Then both Assumptions 2, 3 reduce to computing the prior mass on the $L_{2}$-ball $\bar{S}(r)$, i.e., $\mathbb{P}_{\pi_{0}}(\bar{S}(r))$ for some (sequence of) $r$. In general, Assumption 2 should hold whenever there is sufficient mass around the true parameter $\theta^{*}$, i.e., the prior’s support covers $\theta^{*}$ and Assumption 3 should hold whenever the prior has an exponentially decaying tail.
When the prior has a subexponential tail, Assumption 3 requires the set $\Omega_{n}$ to expand at least exponentially fast with $n$. For instance, it is easy to check that the horseshoe prior satisfies Assumption 3 easily when $\Omega_{n}$ expands with radius $e^{n}$. Following the proof of Lemma 3, one may derive a similar result for $\Omega_{n}$ with radius growing at most $O(e^{n})$ fast, which still hold asymptotically for a wide range of common likelihood functions, e.g., for Gaussian.
The following Lemma is useful when the prior is a standard Gaussian, which also adapts trivially to a spike and slab prior with a Gaussian slab.
@@
Let $\pi_{0}(\cdot)$ and $\bar{S}(r)$ be defined as above, we have two separate lower bounds for the prior mass
$\mathbb{P}_{\pi_{0}}(\bar{S}(r))\geq\begin{dcases}\frac{\exp(-\lambda/2)}{\Gamma(d/2)}\gamma(d/2,r^{2}/2)&\text{ for }r\text{ small}\\ \frac{1}{\Gamma(d/2)}\gamma(d/2,R^{2}/2)&\text{ for }r>\|\theta^{*}\|_{2}.\end{dcases}$
where $\lambda=\|\theta^{*}\|_{2}^{2}$, $R=r-\|\theta^{*}\|_{2}$ and
$\gamma(\alpha,r)=\int_{0}^{r}u^{\alpha-1}\exp(-u)\mathtt{d}u$
is the incomplete gamma function.
@@
$\frac{\gamma(\alpha,r)}{\Gamma(\alpha)}=\,\,$ $\frac{1}{\Gamma(\alpha)}\int_{0}^{r}w^{\alpha-1}e^{-w}\mathtt{d}w$
[…]
where the last step is due to Stirling’s approximation, (see, for instance, (3.9) page 24 of Artin (1964)). Thus $\mathbb{P}_{\pi_{0}}(\bar{S}(r))$ shrinks polynomially with $r$ which is always slower than the exponential shrinkage in Assumption 2 for $n$ large enough.
@@
For $r$ large, to verify Assumption 3, we use the second inequality with the approximation $\Gamma(\alpha)-\gamma(\alpha,{r})\sim r^{\alpha-1}e^{-r}$, or
$\frac{\gamma(\alpha,r)}{\Gamma(\alpha)}\sim 1-\frac{1}{\Gamma(\alpha)}r^{\alpha-1}e^{-r}$
and hence,
$1-\mathbb{P}_{\pi_{0}}(\bar{S}(r))\preceq r^{d-2}\exp(-(r-\|\theta^{*}\|_{2})^{2}/2).$
where $\sim$,$\preceq$ , ⪯ are
  1. 1.
    $f(x)\sim u(x)\text{ if }f/u\rightarrow 1$ as $x\rightarrow L,$ or the leading term of $f(x)$ is $u(x)$ as $x\rightarrow L$;
  2. 2.
    $f_{n}(x_{n})\preceq u_{n}(x_{n})$ if there exists constant $C>0$ independent of $x_{n}$ and $n$ such that $f_{n}(x_{n})\leq Cu_{n}(x_{n})$.
@@ 4 Empirical Assessment of Differential Privacy
It is nontrivial to directly link the $(\epsilon,\delta)$ of the model with the amount of contamination being applied. In other words, it can be hard for the practitioners to directly derive the amount of contamination needed to achieve a certain degree of $(\epsilon,\delta)$ by consulting Theorem 5 along. Thus, this section is devoted to provide guidance on how to empirically estimate $(\epsilon,\delta)$ of the model in practice. Hence, the choice of contamination may be tuned accordingly after the empirical estimation, e.g., increase contamination rate $p_{n}$ if $\epsilon$ is not low enough.
We begin by introducing our pipeline of estimating DP, then followed by some empirical simulations for a few settings of common regression models.
@@ 4.1 Estimation Procedure
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})\leq\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{m_{n}(\Omega,\bm{Z})}{m_{n}(\Omega,\bm{X})}\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{Z}),$
where
$d(\bm{x};\theta):=\frac{k_{p_{n}}(\bm{x};\theta)}{k_{p_{n}}(\bm{x};\theta^{*})}=\frac{(1-p_{n})f(\bm{x};\theta)+p_{n}g(\bm{x})}{(1-p_{n})f(\bm{x};\theta^{*})+p_{n}g(\bm{x})}$
Thus
$\exp(\hat{\epsilon}_{n})\leq\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{m_{n}(\Omega,\bm{Z})}{m_{n}(\Omega,\bm{X})}=\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{\mathbb{E}_{\pi_{n-1}}[d(\bm{z};\theta)]}{\mathbb{E}_{\pi_{n-1}}[d(\bm{x};\theta)]}$
where $\pi_{n-1}$ is the posterior distribution on the common $n-1$ data points shared between $\bm{X}$ and $\bm{Z}$. The first fraction can be dealt with separately for $\bm{x}$ and $\bm{z}$, by solving
$\sup_{\bm{x}\in\mathcal{X}}\sup_{\theta\in A_{n}}d(\bm{x};\theta)\text{ and }\inf_{\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}d(\bm{z};\theta).$ (17)
Since $\pi_{n-1}$ is usually only available up to a constant, we consider solving the Monte Carlo approximated optimization problems
$\sup_{\bm{x}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i}),\quad\inf_{\bm{z}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i})$ (18)
where $\theta_{i}$ are drawn from a proposal distribution, e.g., the Laplace approximation of the posterior distribution (see Section 4.6.8.2, pg. 152-153 of Murphy (2022)),
$\theta_{i}\sim\mathcal{N}(\theta_{\text{MAP}},\bm{H}^{-1})$ (19)
where $\theta_{\text{MAP}}$ is the map estimator of the posterior distribution $\pi_{n-1}$ and $\bm{H}$ is the hessian of $-\log(\pi_{n-1}(\theta|\mathcal{D}))$.
For $\delta$, we will estimate $\mathbb{P}_{\pi_{n}}(A_{n}^{c}|\bm{X})$ as an upperbound for $\delta_{n}$. Again, an analytic solution is usually unavailable, so a numerical scheme is required. Here we recommend using the Laplace approximation again to implement an importance sampling scheme by sampling $\theta_{i}\sim\mathcal{N}(\theta_{\text{MAP}},\bm{H}^{-1})$ and compute
@@
There are two user-defined sequences $\phi_{n}$ and $p_{n}$ that define the radius of the neighbourhood $A_{n}$ and the contamination ratio respectively. In the simulations, we choose
$\phi_{n}:=C_{\phi}^{(1)}\exp(-n^{1/C_{\phi}^{(2)}}),\quad p_{n}:=\exp(-n^{1/C_{p}})$ (21)
where $C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p}$ are constants you may find in the relevant figures/descriptions. Finally, we construct the set $A_{n}$ to be a neighbourhood around $\theta^{*}$, e.g., $L_{\infty}$-ball around $\theta^{*}$. Now, $\epsilon_{n}$ and $\delta_{n}$ may be estimated by Algorithm 1.

### v3

@@ 3.2 DP under Bayesian Sampling
where $\mathbb{P}^{*}_{\theta^{*}}$ is the data generating (outer) measure for the pair of neighbouring datasets of size $n$, $\bm{X}$ and $\bm{Z}$, with entries generated from the true likelihood $k_{p_{n}}(\cdot;\theta^{*})$. In addition, $t_{n}=c_{1}\phi_{n}^{2}/8$, $\psi_{n}$ is defined in Proposition 3, and $c_{5}=\min\{c_{1}\alpha/8,c_{2}\}$. In other words,
$\forall S\in\mathcal{B}(\Omega),\,\,\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp(\epsilon_{n})\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}$
holds with probability tending to one as $n\rightarrow\infty$.
We define in Definition 2 the type of DP proved here, which is adapted from Hall et al. (2013), but in our case, the probability of failing $(\epsilon_{n},\delta_{n})$-DP diminishes exponentially with $n$. The main conclusion of Theorem 4 is that as $n\rightarrow\infty$, we may choose a sequence $p_{n}\rightarrow 0$ such that $\epsilon_{n},\delta_{n}\rightarrow 0$ and for $n$ large enough the failure probability falls into the first case of (14), i.e., $C_{1}\exp(-nc_{5}\phi_{n}^{2})\rightarrow 0$.
@@
For simplicity, we will write $\mathbb{P}$ instead of $\mathbb{P}^{*}_{\theta^{*}}$ when the context is clear. Proposition 1 provides insight into how to upper bound the $\epsilon_{n}$ and $\delta_{n}$. Recall that
@@
Under Assumptions 4 and 5, we can show that $\eta_{n}$ is always finite with $\eta_{n}\rightarrow 1$. We can also characterise its convergence rate (see Lemma 14).
@@
  • In Lemma 12, we prove under Assumption 2 that
@@
    $\mathbb{P}\left(m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})\geq e^{-c_{1}n\phi_{n}^{2}}\right)\leq C^{**}\exp(-c_{2}n\phi_{n}^{2}),$
@@
  • The relevant proofs are contained in the Appendix A, apart from the proof of Lemma 8 in the supplement.
Since we assume that $\bm{Z}$ also follows the same data generating measure, $m_{n}(A_{n}^{c},\bm{Z})$ shares the same upper bound as $m_{n}(A_{n}^{c},\bm{X})$. Now, every term in the inequality has a probabilistic upper bound; combining these bounds completes the proof. ∎
Lemmas 9, 11, and Proposition 3 are intermediate results such that whenever the Assumptions 1 and 2 hold with respect to uncontaminated likelihood $f(\cdot;\theta)$, the same assumptions also hold for the actual contaminated likelihood $k_{p}(\cdot;\theta)$, and in practice, one only needs to check the assumptions for the case $p=0$.
@@
Let $p_{n}\rightarrow 0$ be a sequence of positive real numbers. Let $I_{p_{n}}(\theta)$ denote the Fisher information matrix for a single data point with respect to the density $k_{p_{n}}(\cdot;\theta)$ and $I(\theta)$ denote the Fisher information matrix with respect to $f(\cdot;\theta)$. Then
@@
By picking $p_{n}\rightarrow 0$ such that $\psi_{n}/p_{n}\rightarrow 0$ as $n\rightarrow\infty$, both $\delta_{n}$ and $\epsilon_{n}$ converge to zero. Asymptotically, the algorithm achieves differential privacy for arbitrarily small $(\epsilon,\delta)$. Finally, since $p_{n}\rightarrow 0$, the convergence in the Fisher information follows from Theorem 5.
Our results address the case where samples are drawn exactly from the posterior distribution. For Markov Chain Monte Carlo algorithms (Robert et al., 1999) that do not produce i.i.d. samples, our results indicate the differential privacy level upon reaching stationarity. Furthermore, Minami et al. (2016) can be used to derive a tight bound on the privacy loss incurred when generating an approximate sample, by bounding the total variation distance between the distribution of the approximate sample and the target posterior distribution. There is a growing body of literature on the convergence rates of MCMC algorithms, e.g., Durmus and Moulines (2017), Andrieu et al. (2022), Andrieu et al. (2025), which can be leveraged to complement our analysis.
@@ 3.3 Remarks on Assumptions
Lemmas 8 and 13 jointly provide a probabilistic bound for the posterior mass on $A_{n}^{c}$. This result could alternatively be established via Theorem 8.11 (or Theorem 8.9) of Ghosal and Van der Vaart (2017), under a similar set of assumptions. We adopt the proof techniques from Shen and Wong (1994); Wong and Shen (1995); Shen and Wasserman (2001) to obtain a more concise representation of the bounds along with exponentially shrinking failure probability, in contrast to polynomial decay obtained from the results of Ghosal and Van der Vaart (2017). The assumptions used by Ghosal and Van der Vaart (2017) are, however, arguably less restrictive with respect to the prior and size of $n$, which we discuss in more detail in Appendix D.1.
@@
By Lemma 2, the above holds for most families of likelihood functions $f(\cdot;\theta)$ when $n$ is large enough. However, due to the constant $c$ being quite small in magnitude, see Lemma B.2 in the supplement, one typically requires unrealistically large $n$ for Theorem 4 to hold. Despite having derived its convergence rate in a non-asymptotic fashion, the result is better deemed as an asymptotic one. We will show in the next section using empirical simulations that Theorem 4 provides insight into how $\epsilon_{n}$ and $\delta_{n}$ converge.
Both Assumptions 2 and 3 are conditions on the prior distribution. Let $\bar{S}(r):=\{\theta\in\Omega:\|\theta^{*}-\theta\|_{2}\leq r\}$ denote the $L_{2}$-ball of radius $r$ around $\theta^{*}$. One can show that under mild conditions, the set $S_{0,\alpha}(t_{n})\cap\bar{S}(c_{r})$ contains an $L_{2}$-ball with radius proportional to $t_{n}$ for some constant $c_{r}<\infty$.
@@
Let $S_{0,\alpha}(t_{n}):=\{\theta:\rho_{0,\alpha}(\theta^{*},\theta)\leq t_{n}\}$ defined as in Definition 6. Suppose that there exist $\alpha\in(0,1]$, and $c_{r}<\infty$ such that
$\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\int\left[\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)}\right]^{\alpha+1}\|\tilde{\bm{\nabla}}_{\theta}f(\bm{x};\tilde{\theta})\|_{2}\mathtt{d}x<M_{c_{r}},$
where $M_{c_{r}}$ is a constant depending on $c_{r}$, and $\tilde{\bm{\nabla}}$ is the supremum directional derivative defined in Definition 7. Then $\bar{S}(t_{n}/M_{c_{r}})\subseteq S_{0,\alpha}(t_{n})$.
@@
The condition holds for most likelihood functions, since $\|\tilde{\bm{\nabla}}_{\theta}f(\bm{x};\theta)\|_{2}$ is typically integrable. When $\theta$ does not control the dispersion, $f(\bm{x};\theta^{*})$ and $f(\bm{x};\theta)$ have matching rates of decay and hence the expression is integrable. Otherwise, one can typically choose $c_{r}$ small enough such that the expression remains integrable.
@@
One may use a similar argument to show that the Hellinger distance and the $L_{2}$ distance are locally related (see Appendix C), allowing $\phi_{n}$ and $\psi_{n}$ to be chosen to decay at the same rate. Then, the rate of $p_{n}$ can be chosen freely as long as $\psi_{n}/p_{n}\rightarrow 0$.
Note that we may choose $\Omega_{n}$ in Assumption 3 as an expanding $L_{2}$-ball. Then both Assumptions 2, 3 reduce to computing the prior mass on the $L_{2}$-ball $\bar{S}(r)$, i.e., $\mathbb{P}_{\pi_{0}}(\bar{S}(r))$ for some (sequence of) $r$. In general, Assumption 2 should hold whenever there is sufficient mass around the true parameter $\theta^{*}$, i.e., the prior’s support covers $\theta^{*}$, and Assumption 3 should hold whenever the prior has an exponentially decaying tail.
When the prior has a polynomial tail, Assumption 3 requires the set $\Omega_{n}$ to expand at least exponentially fast with $n$. For instance, the horseshoe prior readily satisfies Assumption 3 when $\Omega_{n}$ expands with radius $e^{n}$. Following the proof of Lemma 2, one can derive a similar result for $\Omega_{n}$ with radius growing at most $O(e^{n})$ fast, which still holds asymptotically for a wide range of common likelihood functions, e.g., Gaussian.
@@
Let $\bar{S}(r)$ be defined as above, and let $\pi_{0}(\cdot)$ be a standard Gaussian prior. We have two separate lower bounds for the prior mass
$\mathbb{P}_{\pi_{0}}(\bar{S}(r))\geq\begin{dcases}\frac{\exp(-\lambda/2)}{\Gamma(d/2)}\gamma(d/2,r^{2}/2),&\text{ for }r\text{ small,}\\ \frac{1}{\Gamma(d/2)}\gamma(d/2,R^{2}/2),&\text{ for }r>\|\theta^{*}\|_{2}.\end{dcases}$
where $\lambda=\|\theta^{*}\|_{2}^{2}$, $R=r-\|\theta^{*}\|_{2}$, and $\gamma(\alpha,r)=\int_{0}^{r}u^{\alpha-1}\exp(-u)\mathtt{d}u$ is the incomplete gamma function.
The above lemma adapts trivially to Gaussian priors with different mean and covariance structures, or to spike and slab priors with Gaussian slabs.
@@
$\frac{\gamma(\alpha,r)}{\Gamma(\alpha)}=\,\,$ $\frac{1}{\Gamma(\alpha)}\int_{0}^{r}w^{\alpha-1}e^{-w}\mathtt{d}w,$
$\geq\,\,$ $\frac{e^{-r}}{\Gamma(\alpha)}\int_{0}^{r}w^{\alpha-1}\mathtt{d}w\geq\frac{1}{2\pi}e^{\alpha}\alpha^{-\alpha-\frac{1}{2}}r^{\alpha}e^{-r},$
where the last step is due to Stirling’s approximation (see, for instance, (3.9) page 24 of Artin (1964)). Thus $\mathbb{P}_{\pi_{0}}(\bar{S}(r))$ shrinks polynomially with $r$, which is always slower than the exponential shrinkage in Assumption 2 for sufficiently large $n$.
@@
For $r$ large, to verify Assumption 3, we use the second inequality with the approximation $\Gamma(\alpha)-\gamma(\alpha,{r})\sim r^{\alpha-1}e^{-r}$, and hence,
$1-\mathbb{P}_{\pi_{0}}(\bar{S}(r))\leq\frac{\Gamma(d/2)-\gamma(d/2,R^{2}/2)}{\Gamma(d/2)}\sim\frac{1}{\Gamma(d/2)}r^{d/2}\exp(-(r-\|\theta^{*}\|_{2})^{2}/2),$
where $\sim$ indicates the leading order term.
@@ 4 Empirical Assessment of Differential Privacy
It is nontrivial to directly link the $(\epsilon,\delta)$ of the model with the amount of contamination being applied. In other words, it can be hard for practitioners to directly derive the amount of contamination needed to achieve a certain degree of $(\epsilon,\delta)$ by consulting Theorem 4 alone. This section is devoted to providing guidance on empirically estimating $(\epsilon,\delta)$ for a given model. The contamination level can then be adjusted based on the empirical estimation, e.g., increase the contamination rate $p_{n}$ if $\epsilon$ is not low enough.
We begin by introducing our pipeline for estimating DP, followed by some empirical simulations for a few settings of common regression models.
@@ 4.1 Estimation Procedure
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})\leq\sup_{\bm{x},\bm{z}\in\mathcal{X}}\left[\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\right]\frac{m_{n}(\Omega,\bm{Z})}{m_{n}(\Omega,\bm{X})}\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{Z}),$
where $m_{n}(\Omega,\bm{X})=\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\pi_{0}(\theta)\mathtt{d}\theta$, and
$d(\bm{x};\theta):=\frac{k_{p_{n}}(\bm{x};\theta)}{k_{p_{n}}(\bm{x};\theta^{*})}=\frac{(1-p_{n})f(\bm{x};\theta)+p_{n}g(\bm{x})}{(1-p_{n})f(\bm{x};\theta^{*})+p_{n}g(\bm{x})}.$
Conditioned on a dataset $\mathcal{D}$ of size n1n-1, we may compute the empirical $\hat{\epsilon}_{n}$ associated with $\mathcal{D}$ by solving
$\exp(\hat{\epsilon}_{n})\geq\sup_{\bm{x},\bm{z}\in\mathcal{X}}\left[\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\right]\frac{m_{n}(\Omega,\mathcal{D}\cup\{\bm{z}\})}{m_{n}(\Omega,\mathcal{D}\cup\{\bm{x}\})}=\sup_{\bm{x},\bm{z}\in\mathcal{X}}\left[\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\right]\frac{\mathbb{E}_{\pi_{n-1}}[d(\bm{z};\theta)]}{\mathbb{E}_{\pi_{n-1}}[d(\bm{x};\theta)]},$
where $\pi_{n-1}$ is the posterior distribution on $\mathcal{D}$. The first fraction can be dealt with separately for $\bm{x}$ and $\bm{z}$, by solving
$\sup_{\bm{x}\in\mathcal{X}}\sup_{\theta\in A_{n}}d(\bm{x};\theta)\text{ and }\inf_{\bm{z}\in\mathcal{X}}\inf_{\theta\in A_{n}}d(\bm{z};\theta).$ (15)
Since $\pi_{n-1}$ is usually only available up to a constant, we may approximate the expectations $\mathbb{E}_{\pi_{n-1}}$ through importance sampling and solve the approximated optimization problems
$\sup_{\bm{x}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i}),\quad\inf_{\bm{z}}\sum_{i=1}^{m}w_{i}d(\bm{z};\theta_{i}),$ (16)
where $(\theta_{i},w_{i})$ are weighted samples for $\pi_{n-1}(\cdot|\mathcal{D})$ with $\theta_{i}$ drawn from some proposal distribution, and $w_{i}$ being the importance weight associated with $\theta_{i}$. For instance, the Laplace approximation of the posterior distribution $\mathcal{N}(\theta_{\text{MAP}},\bm{H}^{-1})$ may be used as the proposal distribution by computing the MAP estimator and the Hessian $\bm{H}$ of $\log(\pi_{n-1})$ (see Section 4.6.8.2, pg. 152-153 of Murphy (2022)).
Remark 13.
There are a few places where the use of $\theta^{*}$ is preferred. An estimate $\hat{\theta}^{*}$ would suffice when $\theta^{*}$ is unknown.
For $\delta$, we will estimate $\mathbb{P}_{\pi_{n}}(A_{n}^{c}|\bm{X})$ as an upper bound for $\delta_{n}$. Again, an analytic solution is usually unavailable, so a numerical scheme is required. Here we recommend using the Laplace approximation again to implement an importance sampling scheme by sampling $\theta_{i}\sim\mathcal{N}(\theta_{\text{MAP}},\bm{H}^{-1})$ and computing importance weight $w_{i}:=\pi_{n}(\theta_{i}|\bm{X})/f_{\mathcal{N}}(\theta_{i};\theta_{\text{MAP}},\bm{H}^{-1})$, where $f_{\mathcal{N}}$ is the Gaussian density function. We may estimate $\hat{\delta}_{n}$ by summing the weights $w_{i}$ of the importance samples $\theta_{i}\in A_{n}^{c}$,
@@
In the proof, we specified the size of the neighbourhood through the sequence $\phi_{n}$. However, this may pose difficulty in controlling $\delta$. For empirical estimation, we instead fix the maximum size of $\delta$ and choose $\phi_{n}$ to be the minimal value for which $\hat{\delta}_{n}<\delta$ in (17). Through $\phi_{n}$, we construct the set $A_{n}$ to be a neighbourhood around $\theta^{*}$, e.g., $L_{\infty}$-ball around $\theta^{*}$. In the simulations, we choose the contamination rate $p_{n}$ through
$p_{n}:=n^{1/8}.$ (18)
Now, $\epsilon_{n}$ may be estimated by Algorithm 1.